Mr. Hernandez had to interview a total of 63
households in his assignment. He has already
finished 28. What percentage of the households
in his assignments has he finished? (Round
your answer to the nearest tenth of a percent.) how i can solve this problem?

Found 2 solutions by peekaboo0501, checkley77:
Answer by peekaboo0501(13)

28 / 63 x 100 = 44.4%

Answer by checkley77(12844)

CREATE A FRACTION WITH THE COMPLETED NUMBER AS THE NUMERATOR & THE TOTAL NUMBER AS THE DENOMINATOR - THEN DIVIDE.
28/63=4/9 OR .444444 OR 44.4% OF THE HOUSES HAVE BEEN FINISHED.